A cylindrical capacitor is constructed using two coaxial cylinders of the same length 10 cm and of radii 2 mm and 4 mm. (a) Calculate the capacitance. (b) Another capacitor of the same length is constructed with cylinders of radii 4 mm and 8 mm. Calculate the capacitance.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. (a) = 2ε base 0L/In(R base 2/R base 1) = e * 3.14 * 8.85 * 10^-2 * 10^-1/In2 [In2 = 0.6932] = 80.17 * 10^-13 ⇒ 8 PF (b) Same as R base 2/R base 1 will be same.